Find the optical density and transmittance of 0.04% -lactose solution, if ⅛ of 

initial light intensity is absorbed in it.


1
Expert's answer
2021-04-22T11:03:05-0400

T"_\lambda" = 1 - "\frac{1}{8}" = "\frac{7}{8}" = 0.875

OD= -log Tλ = -log(0.875) = 0.058